OpenStudy (anonymous):

5,sqrt of 32 + 3, sqrt of 72 =

OpenStudy (anonymous):



OpenStudy (anonymous):

\[38\sqrt{2}\]

OpenStudy (anonymous):

\[2(1+\sqrt[3]{9})\]

OpenStudy (anonymous):

is your question \[\sqrt[5]{32}+\sqrt[3]{72} ??\]or \[5*\sqrt{32}+3*\sqrt{72}\]?

OpenStudy (anonymous):

\[5\sqrt{32} + 3\sqrt{72} = 5\sqrt{16*2} +3\sqrt{36*2} \] \[= 5*4\sqrt 2 + 3*6\sqrt 2 = 20\sqrt 2 + 18\sqrt 2 = 38\sqrt 2 \]

OpenStudy (anonymous):